OpenStudy (anonymous):

what is the nature of the roots of the equation 5n^2=4n+6?

OpenStudy (cwrw238):

first write it as 5n^2 - 4n - 6 = 0 thats in standard form ax^2 + bx + c = 0 the nature of its roots is found by working out the value of the discriminant which is b^2 - 4ac where a,b and c are the values in the standard form

OpenStudy (cwrw238):

so if you compare 5n^2 - 4n - 6 = 0 with the standard form a = 5 , b = -4 and c = -6 right? do you follow that? ( ignore difference in the variables x and n )

OpenStudy (anonymous):

yes i follow :)

OpenStudy (cwrw238):

so you have b^2 - 4ac = (-4)^2 - 4*5*-6 if this is > 0 there are 2different real roots i this = 0 there are 2 duplicate real roots if < 0 there are no real roots

OpenStudy (anonymous):

yes I have that. should my answer be 136?

OpenStudy (cwrw238):

yep - so how many roots are there and are they real?

OpenStudy (anonymous):

i'm not sure about roots -_-

OpenStudy (cwrw238):

the reason we use the discriminant is because its part of the formula for roots of a quadratic equation and its square roots is taken x = [-b +/- sqrt(b^2-4ac)] / 2a

OpenStudy (cwrw238):

the roots of an equation are also called the zeros

OpenStudy (cwrw238):

the answer here is 2 real zeros because the discriminant is = 136 which is > 0

OpenStudy (anonymous):

Ohh okay! That's what I wasn't understanding, sorry. I'm trying to study for an exam that's tomorrow